OpenStudy (anonymous):

Find the slope of the tangent line to the curve 4x^2-2xy+2y^3=94 at the point (4,3) .

OpenStudy (abb0t):

Find the derivative and then plug in x = 4 and y = 3. That will be your slope of the tangent line to the given curve.

OpenStudy (abb0t):

This involved implicit differentiation.

OpenStudy (anonymous):

i found the derivative, but it wasn't right still?

OpenStudy (dumbcow):

after differentiating \[8x-2y-2x \frac{dy}{dx}+6y^{2} \frac{dy}{dx} = 94\] solving \[\frac{dy}{dx} = \frac{2y-8x+94}{6y^{2}-2x}\] plug in point (4,3) \[\frac{dy}{dx} = \frac{34}{23}\]